OpenStudy (anonymous):

Solve the system for  y.  x −  y  =  6  5 x + 4 y  =  − 24

OpenStudy (zale101):

Use either one of the equation and solve for x the equation you got for x, plug it into the other x in the second equation.

OpenStudy (anonymous):

I tried isolating my y for the 5x+4y=-24 and it doesn't work! I don't get it ...

OpenStudy (zale101):

you'll only be left with y variable to solve for it.

OpenStudy (anonymous):

for the second one x= −4 5 y+ −24 5

OpenStudy (zale101):

If you isolate y first, then you'll have to solve for x. If you do this method, you'll have to do an extra step where you have to plug what you got for x back into either one of the equations to get y.

OpenStudy (anonymous):

how is the problem supposed to look like when first starting it?

OpenStudy (zale101):

 x −  y  =  6  5 x + 4 y  =  − 24 i'll choose the first equation and solve for x \(x-y=6\) \(x=6+y\)

OpenStudy (zale101):

Makes sense?